OpenStudy (anonymous):

need help with 6 question I need to get done please help me

OpenStudy (anonymous):

21. ABCD is a rectangle. find the length of BD, if AC=5x+1 and BD=6x-2

OpenStudy (juicstice):

just set the problems equal to each other 5x+1=6x-2 then solve for x

OpenStudy (anonymous):

then

OpenStudy (juicstice):

then you just plug x into the equation for BD and you have it

OpenStudy (anonymous):

u need help

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay with

OpenStudy (anonymous):

with this qustion

OpenStudy (juicstice):

do you not understand my explanation?

OpenStudy (anonymous):

I not im trying I just not get math at all

OpenStudy (juicstice):

okay well its really simple I'll do my best at drawing it

OpenStudy (anonymous):

sam i inbox u

OpenStudy (anonymous):

ok thx u

OpenStudy (anonymous):

do u know me

OpenStudy (anonymous):

you got a picture? this doesn't seem right to me

OpenStudy (anonymous):

there is no picture

OpenStudy (anonymous):

ok can you solve \[5x+1=6x-2\] for \(x\) ?

OpenStudy (juicstice):



OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sam inbox me

OpenStudy (juicstice):

did you need help with any more?